sweetJ12:

the length of the two different cubes are 36 cm and 45 cm what is the ratio of the volume of the smaller to the volume of the larger (in simplest form) 16/25 64/125 4/5 12/15

Vocaloid:

volume of a cube = s^3 where s = the side length use this formula to calculate: smaller cube volume/larger cube volume, then reduce the fraction to lowest terms let me know if you are getting stuck or just want to check your solution

sweetJ12:

so it would be 4/5?

Vocaloid:

I got something a little different small volume/large volume = 36^3 / 45^3 = ?

Vocaloid:

^3 means cubed 36^3 / 45^3 = [36*36*36] / [45*45*45] = ?